40. The labour cost to construct the first unit of a special piece of equipment was $126,000. If labour costs are subject to a 90% learning curve, what will be the cost to build the second unit? Use the following values as needed to respond In(95%) / ln(2)s-0074 In(90%) /ln(2) =-0.152 In(8596) / ln(2) =-0.234 In(80%) / In(2) =-0.322 n(75%)/In(2) =-0.415 In(70%)/In(2) -0.515 a) $126,000 b) $113,000 c) $100,800 d) $113,400 Answer: d
41 A firm will build 8 units of a product. All costs are subject to an 85% learning curve. The first unit cost $75,000 to build. If the firm values inventory and cost of goods sold at the average cost for all units in the production run, what will be the cost to produce the second unit? Use the following values as needed to respond: In(95%) / In(2)--0074 In(90%) /ln(2) #-0.152 In(8596) / In(2)--0234 In(80%) / ln(2) -0.322 In(75%)/In(2) -0.415 In(70%)/In(2)s-0.515 a) $63,750 b) $69,375 c) $75,000 d) Some other amount Answer: b Difficulty: Easy Learning Objective: Describe cost estimation techniques CPA Management Aspounting/Problem-Solving and Decision-Making Bloomcode: Application 42 Assume an 80% learning curve and that the first unit takes 6 hours. How long does it take to produce the second unit? Use the following values as needed to respond: In(9596) / In(2) -0.074 n(90%)/In(2)--0.152 In(8596) / ln(2) =-0.234 In(80%) /ln(2) -0.322 In(75%) /ln(2) #-0415 In(70%) /ln(2)--0515 a) 4.1 hours 3.6 hours c) 4.8 hours d) Some other amount Answer c
43. Managers at Bobs Custom Saddle Manufacturers believe that the learning rate for new employees is about 90%. If it takes a new employee 10 hours to make the first saddle, estimate the cumulative average time per saddle for a new employee to make four saddles. Use the following values as needed to respond: In(95%) / ln(2) -s-0074 In(90%) /ln(2) =-0.152 In(85%) / ln(2) ::-0.234 In(80%) / ln(2)--0322 In(75%)/In(2) -0.415 In(70%)/ln(2)--0.515 a) 8.1 hours per saddle b) 5.3 hours per saddle c) 7.6 hours per saddle d) Cannot be determined Answer: a Difficulty: Easy Learning Objective: Describe cost estimation techniques. CPA: Management Accounting/Problem-Solving and Decision-Making Bloomcode: Application 44. The cumulative average-time learning curve can be represented mathematically as: a) In(percent learning)/ In(2) Answer: b Difficulty: Easy Learning Objective: Describe cost estimation techniques CPA: Management Accounting Bloomcode: Knowledge
46. Tom and Mike are installing new flooring in their home. The house has 5 rooms of approximately equal size where the flooring will be replaced. Tom and Mike spent 3 hours removing the old flooring from the first room. If Tom and Mike are operating with an 85% learning curve, what is the estimated average time to remove the flooring from all 5 rooms? Note: In(8596) / ln(2) -0.234 a) 2.05 hours total b) 2.05 hours per room c) 2.55 hours total d) 2.55 hours per room Answer: b Difficulty: Easy Learning Objective: Describe cost estimation techniques. CPA: Management Accounting/Problem-Solving and Decision-Making Bloomcode: Application
56. Bava Company wants to develop a cost function for its maintenance costs to estimate such costs for the coming year. The following data are available Maintenance Month June 4,000 6,500 7,000 5,500 $ 900 1,325 1,500 1,150 July September Using the high-low method, what is the variable maintenance cost per direct labour hour? a) $1.00 b) $0.10 c) $0.20 d) $1.50 Answerc Difficulty: Easy Learning Objective: Describe cost estimation techniques. CPA: Management Accounting/Problem-Solving and Decision-Making Bloomcode: Application 57. Bava Company wants to develop a cost function for its maintenance costs to estimate such costs for the coming year. The following data are available Direct Maintenance Month June 4,000 $ 900
July August September 6,500 7,000 5,500 1,325 1,500 1,150 Using the high-low method, what is the fixed maintenance cost? a) $500 b) $300 c) $200 d) $100 Answer: d Difficulty: Easy Learning Objective: Describe cost estimation techniques. CPA: Management Accounting/Problem-Solving and Decision-Making 58. Bava Company wants to develop a cost function for its maintenance costs to estimate such costs for the coming year. The following data are available: Direct Maintenance June July August September 4,000 6,500 7,000 5,500 $900 1,325 1,500 1,150 Using the high-low method, what is the cost function for maintenance costs? a) $500+$1.00 per direct labour hour b) $300 +$1.50 per direct labour hour )$100+$O.20 per direct labour hour d) $200+$O.10 per direct labour hour Answer: c Difficulty: Easy Learning Objective: Describe cost estimation techniques. CPA: Management Accounting/Problem-Solving and Decision-Making Bloomcode: Application
Answer #1

40. Average Labor Cost to manufacture two units (y)

y = a * x^b

y = 126000 * 2 ^ - 0.152

y = 126000 * 0.9

Cost of Second Unit = $113400 Option d

41.

Average Learning Curve for the second unit =100% - [ (100 % - 85%) / 2 ] = $92.50%

Average Cost to manufacture two units (y)

y = a * x^b

y = 75000 * 2 ^ - 0.112

y = 75000 * 0.925

Cost of Second Unit = $69375 Option b

42. Time for second unit = First Unit time * Total Units ^ {Ln(80%)/ Ln(2)]

Time for second unit = 6 * 0.80

Time for second unit = 4.80 Hours Option C

43. Estimated cumulative average time = First Unit time * Total Units ^ ln(90%)/ln(2)

Estimated cumulative average time = 10 * 4 ^ -0.152

Estimated cumulative average time = 10 * 0.81

Estimated cumulative average time = 8.1 Hours Option A

As per HOMEWORKLIB RULES experts are required to answer only first 4 questions when multiple questions were asked. please ask remaining questions separately. Thanks for understanding